I need help with #4, 6, 7, 8, and 9

Bertha is a 65 year old Caucasian female with an admission height of 5’4” and weight of 135 lb. was admitted with a severe episode of CVA. Considering that her recent CVA affected her speech, she was unable to communicate her weight and diet intake history to the admitting nurse. Within 24 hours, her attending primary care physician placed orders for a complete metabolic panel, Mechanical Soft/Cardiac Diet, speech consult, and nutrition consult. Upon completion of the speech pathology assessment, she documented that the patient is at high risk for aspiration and inappropriate for oral feeding at this time. During the initial nutrition assessment, the RD interviews the patient’s daughter to obtain information on her weight history and adequacy of oral intake. Daughter reports that patient has a normal weight range of 130-135 lb., has a good appetite, but doesn’t closely monitor her intake of sodium and refined sweets as often as she should. Furthermore, the MD and speech pathologist estimate that Bertha is expected to need nutrition support for 5 weeks while undergoing rehabilitation.

Past Medical History: HTN, hysterectomy, Depression, GERD

Medications: Prilosec, Lisinopril, Zoloft

Pertinent Admission Labs:

Parameter

Lab Value

BUN

43

Creatinine

0.8

Sodium

153

Potassium

4.3

Albumin

3.0

Magnesium

2.8

Phosphorus

3.5

Glucose

108

  1. As the clinical nutritionist, which alternative feeding modality would you recommend and why?
  2. Would you suggest inserting a long or short-term feeding device? Which device of the chosen feeding modality (enteral or parenteral) would be appropriate for the circumstances of this patient and why?
  3. What issues are this patient’s current lab values suggesting?
  4. What are her estimated calorie, protein, and fluid requirements? Please show all calculations.
  5. Construct two appropriate PES statements for this case.
  6. Intervention: What feeding rate/hour would you recommend daily? If this patient is on enteral support, how many ml of free water would you recommend? If this patient is on parenteral support, what is this patient’s GIR (Glucose Infusion Rate) and what would be the appropriate dose of dextrose, AA, and lipid content? Please show all calculations.
  7. What classification of risk is this patient (low, moderate, high) and within how many days will you follow up on patient progress?
  8. What members of the healthcare team will you communicate with for coordination of care? Explain why.
  9. What would be your Monitoring and Evaluation plan for this patient case?
Answer #1

1. As a nutritionalist ,I would prefer for parental route as an alternative method of feeding because this will help to give both diet and also the medications .

2. Parenteral route would be appropriate for the patient because it would help the healthcare workers to give both medications which needs to be given in infusion and also diet in the form of TPN can be given without any extra lines

3. The issues in the patients lab values are :-

- increased BUN

- increased sodium levels

4. Weight of the patient is135 lb = 135 × 0.45 kg

= 60.75kg

Find your body weight in kilograms (if you live in the US, just divide your weight in pounds by 2.2 to get your weight in kilograms).

= 60.75kg

Step 2: Multiply your weight in kilograms by 0.9 if you are a woman

60.75× 0.9= 54.675

Step 3: Multiply by 24

= 54.675× 24 = 1312.2

Step 4: Multiply by your “Lean Factor”

= 1312.2× 0.85( lean factor for this patient )

= 1115.37

Step 5 : Multiply by your activity modifier

= 1115.37 × 1.55 (activity modifier -light )

= 1728.8 calories /day

#. Protein intake = 0.8 ( minimum protein per day ) × 60.75

= 48.6 gm protein per day is her requirement

# Fluid requirement :

For > 20 kg = 1500 mL + [weight (kg) x 20 ml/kg/day]

= 1500ml + [ 60.75× 20]

= 1500ml + 1215

= 2715 ml /day